The following table contains the probability distribution for X = the number of traffic accidents reported in a day in Corvallis, Oregon.

X

1

5

9

12

15

41

P(X)

0.05

0.10

0.35

0.35

0.10

0.05



The variance of the number of accidents is ________.

(Round your answer to three decimal places.)
Answer #1
x P(x) xP(x) x2P(x)
1 0.05 0.050 0.050
5 0.1 0.500 2.500
9 0.35 3.150 28.350
12 0.35 4.200 50.400
15 0.1 1.500 22.500
41 0.05 2.050 84.050
total 11.450 187.850
E(x) =μ= ΣxP(x) = 11.4500
E(x2) = Σx2P(x) = 187.8500
Var(x)=σ2 = E(x2)-(E(x))2= 56.748

variance of the number of accidents is =56.748